In the town of bayland their are 2 times as many spanish-speaking people as their are french-speaking people and 5 times as many english-speaking people as thier are spanish and french speaking. If the town of bayland has a total population of 2250. How many english speaking people are there ?
Answer by ptaylor(2198)


Let x=number of French-speaking people
Then 2x=number of Spanish-speaking people
And 5(x+2x)=15x=number of English-speaking people
Now we are told that:
x+2x+15x=2250 or
18x=2250 divide each side by 18
x=125-------------------------------number of french-speaking people
2x=250=-----------------------------number of Spanish-speaking people
14x=15*125=1875----------------------number of English-speaking people
Ck
1875+250+125=2250
2250=2250
Hope this helps---ptaylor